# Building Bridges, Not Walls - Advancing Interpretability by Unifying Feature, Data, and Model Component Attribution

### v1

@@ Abstract
The increasing complexity of AI systems has made understanding their behavior a critical challenge. Numerous methods have been developed to attribute model behavior to three key aspects: input features, training data, and internal model components. However, these attribution methods are studied and applied rather independently, resulting in a fragmented landscape of approaches and terminology. This position paper argues that feature, data, and component attribution methods share fundamental similarities, and bridging them can benefit interpretability research. We conduct a detailed analysis of successful methods across three domains and present a unified view to demonstrate that these seemingly distinct methods employ similar approaches, such as perturbations, gradients, and linear approximations, differing primarily in their perspectives rather than core techniques. Our unified perspective enhances understanding of existing attribution methods, identifies shared concepts and challenges, makes this field more accessible to newcomers, and highlights new directions not only for attribution and interpretability but also for broader AI research, including model editing, steering, and regulation.
@@ 1 Introduction
As AI systems grow increasingly complex, understanding their behavior remains a critical challenge (Arrieta et al., 2020; Longo et al., 2024). Researchers have developed methods to explain AI systems by attributing their behavior to three distinct aspects: input features, training data, and internal model components. Feature attribution methods identify influence of input features at test time, revealing which aspects of the input drive the model’s output (Zeiler and Fergus, 2014; Ribeiro et al., 2016; Lundberg and Lee, 2017; Smilkov et al., 2017). Data attribution analyzes how training data shape model behavior during the training phase (Koh and Liang, 2017; Ghorbani and Zou, 2019; Ilyas et al., 2022). Component attribution examines the internal workings of the model by analyzing how specific components, such as neurons or layers in a neural network (NN), affect model behavior (Vig et al., 2020; Meng et al., 2022; Nanda, 2023; Shah et al., 2024). While numerous methods have been developed for each of these types of attribution, and some survey papers have been published (Guidotti et al., 2018; Covert et al., 2021; Wang et al., 2024; Hammoudeh and Lowd, 2024; Bereska and Gavves, 2024), they have been studied and used rather independently by different communities, creating a fragmented landscape of methods and terminology.
Our position is that feature, data, and component attribution methods can be bridged to advance not only interpretability research, by stimulating cross-domain knowledge transfer, but also broader AI research, including model editing, steering, and regulation. We show that these three domains employ common approaches and they differ primarily in perspective rather than core techniques. In the following sections, we first formalize a unified attribution problem that encompasses all three attribution types to show these seemingly distinct domains fall under the same framework (§2). We then examine each domain’s evolution and analyze its successful methods, revealing how these methods are connected through shared techniques and concepts, including perturbations, gradients, and linear approximations (§3, §4, §5). Building on this analysis, we present a unified view and illustrate shared concepts (§6.1), identify common challenges (§6.2), and highlight how this unified perspective facilitates cross-domain knowledge transfer for new research development in interpretability (§6.3), and broader AI research (§6.4). We also discuss alternative views of attribution unification (§7). In summary, we believe that this unified view enhances our understanding of attribution methods, bridges the current fragmented landscape, makes the field more accessible to newcomers, and provides new insights and research directions.
@@ 2 The Attribution Problem
Figure 1: The three types of attribution regarding input features, training data, and internal model components. Each type provides complementary insight into model behavior by analyzing different elements of the AI system with similar methods.
@@ 3.2 Gradient-Based Feature Attribution
Gradient-based feature attribution emerged from computer vision applications, where it gained widespread adoption for generating attribution scores as image saliency maps (Simonyan et al., 2013), also known as sensitivity maps (Smilkov et al., 2017). The “vanilla gradients” method uses the gradients of the output class (log)probability with respect to input pixels as attribution scores (Simonyan et al., 2013). Since then, researchers have proposed numerous enhanced gradient-based methods. For example, Gradients $\times$ Input (Shrikumar et al., 2017) multiplies gradients with input values, Integrated Gradients (Sundararajan et al., 2017) accumulates gradients along a path from a baseline to the actual input, and Integrated Hessians (Janizek et al., 2021) further extends the analysis to feature interactions by computing the Hessian matrix. These methods leverage different gradient formulations to provide more accurate and stable attribution scores. A notable advancement is SmoothGrad (Smilkov et al., 2017), which generates multiple copies of the input with added Gaussian noise and computes sensitivity maps for each noisy sample. By averaging these maps, SmoothGrad reduces noise while preserving salient features that consistently influence model outputs.
@@ 5.3 Linear Approximation for Component Attribution
Given the rapid increase in model size and the combinatorial nature of searching for effective components, component attribution also employs linear approximations like LIME and Datamodels. COAR Shah et al. (2024) attempts to decompose model behavior in terms of various model components by predicting the counterfactual impact of ablating each component, similar to many forms of causal mediation analysis. Given the computational complexity of this problem, they employ linear approximations by assigning scores to each component of a model and estimating the counterfactual effect of removing sets of components by simply summing their corresponding scores. Thus, the complexity of relationships between components is abstracted away through the linear approximation.
@@ 6 Position and Contributions
Feature, data, and component attribution methods have largely been studied as separate domains, resulting in the parallel development of similar methods from different communities with distinct terminologies. We argue that these methods can be unified into a holistic view. Having demonstrated their methodological similarities of three domains, we now summarize their common concepts and challenges and identify promising research directions through cross-domain knowledge transfer. We believe that our unified view will bridge the current fragmented landscape, make the field more accessible to newcomers, and help advance research in interpretability and beyond.
@@ 6.1 Common Concepts of Attribution Methods
As we discussed in the previous sections, attribution methods across features, data, and components can be categorized into: perturbation-based, gradient-based, and linear approximation. We provide a detailed discussion of these categories in Appendix B with Table 2 summarizing all the methods we discussed. Beyond algorithmic similarities, conceptual ideas also transfer across areas of attribution. One example is the deliberate introduction of randomness and smoothing to enhance attribution robustness. This idea has proven effective across domains through random noisy samples used by SmoothGrad for feature attribution, attribution results ensembled over multiple retrainings by TRAK for data attribution, and aggregated results from multiple corrupted runs in causal mediation analysis for component attribution. Another example is tracking and aggregating results along paths, as Integrated Gradients along paths from base input to target input in feature attribution, TracIn tracing training paths to reveal dynamic data influences in data attribution, and path patching tracking component effects along residual stream paths in component attribution. These cross-domain concepts highlight the fundamental connections of attribution methods.
@@ 6.2 Common Challenges of Attribution Methods
Consistency Challenges refer to the variability in attribution results across multiple runs of the same method with different random seeds, making it challenging to establish stable interpretations and evaluations. This challenge is prevalent across all attribution types due to multiple sources of randomness in their common techniques, including sampling, learning processes with stochastic optimization, and also non-trivial hyperparameters. While some gradient-based methods can produce consistent results in a single run when they do not involve sampling or approximations for computationally intensive operations, the consistency between different gradient-based methods varies considerably, which is a problem for all three method categories and leads to the following evaluation challenges.
Evaluation Challenges arise in all three domains due to multiple factors. Inconsistent results make it difficult to reliably compare methods and determine their relative accuracy. This challenge is particularly evident in gradient-based feature attribution methods. As demonstrated by Adebayo et al. (2018), they can produce contradictory attribution results and sometimes perform no better than random baselines, making them difficult to evaluate fairly. For existing evaluation metrics, counterfactual evaluation could provide more rigorous validation, but computational constraints often make this approach impractical. Task-specific evaluations offer easier alternatives but frequently lack generalizability across different contexts. Human evaluation, despite being considered the gold standard, faces scalability issues and potential biases. The diversity of evaluation metrics and their varying definitions of importance make the evaluation more challenging. An attribution result may perform well by one metric but poorly according to another. These challenges emphasize the pressing need for developing more reliable and practical attribution evaluation metrics.
6.3 Cross-Domain Innovation: Leveraging Insights Across Attribution Methods
The connections among feature, data, and component attributions discussed in the sections above suggest multiple promising directions for future research. One research direction is to leverage insights from one area to develop methods for another. This can be directly identified as filling in the empty cells in Table 2. For example, while the Shapley value has been successfully applied across all three attribution domains, many other game-theoretic notions have only been used in feature attribution and not for data and component attribution. In addition, some advanced gradient techniques are common for feature and data attribution methods, but not for component attribution. The Hessian matrix, for example, has been used to obtain second-order information in Integrated Hessians for feature attribution and extensively in all IF-related methods for data attribution, and can be explored for component attribution.
@@ 6.3 Cross-Domain Innovation: Leveraging Insights Across Attribution Methods
6.4 Connections to Other Areas of AI Research
@@ 7 Alternative Views
One alternative view suggests that feature, data, and component attribution represent fundamentally distinct research directions, each with its own terminology and research community. Despite using similar methods, these fields differ in their motivations and methodological approaches. The term “attribution” carries unique meanings in each domain. Feature attribution has historically been synonymous with “explainable AI,” data attribution was part of “data-centric AI” focusing on concrete applications like data selection, and component attribution has emerged recently as one branch of “mechanistic interpretability” to reverse-engineer algorithms in foundation models. This rather independent evolution has produced distinct best practices within each domain (Guidotti et al., 2018; Hammoudeh and Lowd, 2024; Bereska and Gavves, 2024) and separate communities that may seem to resist unification. Nevertheless, we argue that meaningful unification across these fields is possible and will help advance interpretability research.
Researchers have also proposed various taxonomies within each attribution domain. Feature attribution methods, for example, were classified as model-agnostic or model-specific approaches (Arrieta et al., 2020). Model-agnostic methods treat models as black boxes and generate attributions based solely on input-output relationships, while model-specific methods rely on known model architectures. Although these categorizations aid method selection, they have limited generalizability across domains. Component attribution methods, for instance, are inherently model-specific since they analyze defined model components. Our unification approach extends beyond domain-specific frameworks to take a broader, cross-domain perspective. This comprehensive view uncovers connections that remain obscured when examining methods in isolation. Consider gradient-based approaches: while all three domains employ gradients, they compute them for different purposes with different objectives. Thus, while domain-specific categorizations offer valuable insights, our cross-domain framework provides a holistic understanding of the attribution landscape.
@@ 8 Conclusion
In this paper, we have presented a unified view on three traditionally separate attribution methods for interpretability: feature attribution, data attribution, and component attribution. While these methods have evolved independently to explain different aspects of AI systems we demonstrate that they share fundamental technical building blocks such as perturbations, gradients, and linear approximations. This unified view not only bridges the fragmented landscape of attribution methods but also identifies common challenges and opportunities for cross-domain knowledge transfer, ultimately working toward more comprehensive approaches to advancing AI interpretability research.
@@ Appendix B Summary of Methods
In total, there are thousands of attribution methods from all three domains proposed in the literature making a comprehensive summary infeasible. In Table 2, we summarize the attribution methods discussed in this paper, which we believe are the representative ones and align with the unified view we presented. Most of the empty cells in Table 2 (labeled as “–”) represent methods that we believe are promising but have not yet been explored in the literature, as discussed in § 6.3. These represent research ideas that have been verified in one attribution domain but remain unexplored in others. The exception is mask learning for data attribution, which we consider less promising because learning a high-dimensional mask of size $n$ jointly with the model would be infeasible when the model has not been trained.
Table 2: Summary of feature, data, and component attribution methods from all categories.
Method Feature Attribution Data Attribution Component Attribution
Perturb Direct Occlusions [Zeiler and Fergus, 2014] LOO [Cook and Weisberg, 1982] Causal Tracing [Meng et al., 2022]
RISE [Petsiuk, 2018] Path Patching [Wang et al., 2022]
Vig et al. [2020]
Bau et al. [2020]
ACDC [Conmy et al., 2023]
Game-Theoretic SHAP [Lundberg and Lee, 2017] Data Shapley [Ghorbani and Zou, 2019] Neuron Shapley [Ghorbani and Zou, 2020]
(Shapley) TMC Shapley [Ghorbani and Zou, 2019]
KNN Shapley [Jia et al., 2019]
Beta Shapley [Kwon and Zou, 2022]
Game-Theoretic STII [Dhamdhere et al., 2019] Data Banzhaf [Wang and Jia, 2023]
(Others) BII [Patel et al., 2021]
Core Value [Yan and Procaccia, 2021]
Myerson Value [Chen et al., 2018b]
HN Value [Zhang et al., 2022]
Mask Learning Dabkowski and Gal [2017] Csordás et al. [2020]
L2X [Chen et al., 2018a] Subnetwork Pruning [Cao et al., 2021]
Gradient First-Order Vanilla Gradients [Simonyan et al., 2013] GradDot/GradCos [Pruthi et al., 2020] Attribution Patching [Nanda, 2023]
Gradient $\times$ Input [Shrikumar et al., 2017] EAP [Syed et al., 2023]
SmoothGrad [Smilkov et al., 2017]
GBP [Springenberg et al., 2014]
Grad-CAM [Selvaraju et al., 2016]
Second-Order Integrated Hessian [Janizek et al., 2021] IF [Koh and Liang, 2017]
(Hessian/IF) FastIF [Guo et al., 2021]
Arnoldi IF [Schioppa et al., 2022]
EK-FAC [Grosse et al., 2023]
RelateIF [Barshan et al., 2020]
Tracing Path Integrated Grad [Sundararajan et al., 2017] TracIn [Pruthi et al., 2020] Attribution Path Patching [Nanda, 2023]
SGD-Influence [Hara et al., 2019]
SOURCE [Bae et al., 2024]
Linear LIME [Ribeiro et al., 2016] Datamodels [Ilyas et al., 2022] COAR [Shah et al., 2024]
C-LIME [Agarwal et al., 2021] TRAK [Park et al., 2023]

### v2

# Building Bridges, Not Walls: Advancing Interpretability by Unifying Feature, Data, and Model Component Attribution

@@ Abstract
The increasing complexity of AI systems has made understanding their behavior a critical challenge. Numerous methods have been developed to attribute model behavior to three key aspects: input features, training data, and internal model components. However, these attribution methods are studied and applied rather independently, resulting in a fragmented landscape of approaches and terminology. This position paper argues that feature, data, and component attribution methods share fundamental similarities, and bridging them can benefit interpretability research. We conduct a detailed analysis of successful methods of these three attribution aspects and present a unified view to demonstrate that these seemingly distinct methods employ similar approaches, such as perturbations, gradients, and linear approximations, differing primarily in their perspectives rather than core techniques. Our unified perspective enhances understanding of existing attribution methods, identifies shared concepts and challenges, makes this field more accessible to newcomers, and highlights new directions not only for attribution and interpretability but also for broader AI research, including model editing, steering, and regulation.
@@ 1 Introduction
As AI systems grow increasingly complex, understanding their behavior remains a critical challenge (Arrieta et al., 2020; Longo et al., 2024). Researchers have developed methods to explain AI systems by attributing their behavior to three distinct aspects: input features, training data, and internal model components. Feature attribution methods identify influence of input features at test time, revealing which aspects of the input drive the model’s output (Zeiler and Fergus, 2014; Ribeiro et al., 2016; Horel and Giesecke, 2020, 2022; Lundberg and Lee, 2017; Smilkov et al., 2017). Data attribution analyzes how training data shape model behavior during the training phase (Koh and Liang, 2017; Ghorbani and Zou, 2019; Ilyas et al., 2022). Component attribution examines the internal workings of the model by analyzing how specific components, such as neurons or layers in a neural network (NN), affect model behavior (Vig et al., 2020; Meng et al., 2022; Nanda, 2023; Shah et al., 2024). While numerous attribution methods have been developed for each of these three aspects, and some survey papers have been published (Guidotti et al., 2018; Covert et al., 2021; Wang et al., 2024; Hammoudeh and Lowd, 2024; Bereska and Gavves, 2024), they have been studied and used rather independently by different communities, creating a fragmented landscape of methods and terminology for similar ideas (Saphra and Wiegreffe, 2024).
Our position is that feature, data, and component attribution methods can be bridged to advance not only interpretability research, by stimulating cross-aspect knowledge transfer, but also broader AI research, including model editing, steering, and regulation. We show that these three types of attribution employ common methods and they differ primarily in perspective rather than core techniques. In the following sections, we first formalize a unified attribution problem that encompasses all three aspects to show these seemingly distinct approaches fall under the same framework (§2). We then examine the evolution of each attribution aspect and analyze its successful methods, revealing how these methods are connected through shared techniques and concepts, including perturbations, gradients, and linear approximations (§3, §4, §5). We summarize the representative attribution methods in these categories in Table 1. Building on this analysis, we present a unified view and illustrate shared concepts (§6.1), identify common challenges (§6.2), and highlight how this unified perspective facilitates cross-aspect knowledge transfer for new research development in interpretability (§6.3), and broader AI research (§6.4). We also discuss alternative views of attribution unification (§7). In summary, we believe that this unified view enhances our understanding of attribution methods, bridges the current fragmented landscape, makes the field more accessible to newcomers, and provides new insights and research directions.
@@ 2 The Attribution Problem
Figure 1: The three types of attribution regarding input features, training data, and internal model components. Each type provides complementary insight into model behavior by analyzing different aspects of the AI system with similar methods.
Table 1: A summary of representative feature, data, and component attribution methods classified into three methodological categories demonstrating our unified view.
Method Feature Attribution Data Attribution Component Attribution
Perturb Direct Occlusions (Zeiler and Fergus, 2014) LOO (Cook and Weisberg, 1982) Causal Tracing (Meng et al., 2022)
RISE (Petsiuk, 2018) Path Patching (Wang et al., 2022)
Vig et al. (2020)
Bau et al. (2020)
ACDC (Conmy et al., 2023)
Game-Theoretic SHAP (Lundberg and Lee, 2017) Data Shapley (Ghorbani and Zou, 2019) Neuron Shapley (Ghorbani and Zou, 2020)
(Shapley) TMC Shapley (Ghorbani and Zou, 2019)
KNN Shapley (Jia et al., 2019)
Beta Shapley (Kwon and Zou, 2022)
Game-Theoretic STII (Dhamdhere et al., 2019) Data Banzhaf (Wang and Jia, 2023)
(Others) BII (Patel et al., 2021)
Core Value (Yan and Procaccia, 2021)
Myerson Value (Chen et al., 2018b)
HN Value (Zhang et al., 2022)
Mask Learning Dabkowski and Gal (2017) Csordás et al. (2020)
L2X (Chen et al., 2018a) Subnetwork Pruning (Cao et al., 2021)
Gradient First-Order Vanilla Gradients (Simonyan et al., 2013) GradDot/GradCos (Pruthi et al., 2020) Attribution Patching (Nanda, 2023)
Gradient $\times$ Input (Shrikumar et al., 2017) EAP (Syed et al., 2023)
SmoothGrad (Smilkov et al., 2017)
GBP (Springenberg et al., 2014)
Grad-CAM (Selvaraju et al., 2016)
Second-Order Integrated Hessian (Janizek et al., 2021) IF (Koh and Liang, 2017)
(Hessian/IF) FastIF (Guo et al., 2021)
Arnoldi IF (Schioppa et al., 2022)
EK-FAC (Grosse et al., 2023)
RelateIF (Barshan et al., 2020)
Tracing Path Integrated Grad (Sundararajan et al., 2017) TracIn (Pruthi et al., 2020) Attribution Path Patching (Nanda, 2023)
SGD-Influence (Hara et al., 2019)
SOURCE (Bae et al., 2024)
Linear LIME (Ribeiro et al., 2016) Datamodels (Ilyas et al., 2022) COAR (Shah et al., 2024)
C-LIME (Agarwal et al., 2021) TRAK (Park et al., 2023)
@@ 3.2 Gradient-Based Feature Attribution
Gradient-based feature attribution emerged from computer vision, where it gained widespread adoption for generating attribution scores as image saliency maps (Simonyan et al., 2013), also known as sensitivity maps (Smilkov et al., 2017). The “vanilla gradients” method uses the gradients of the output class (log)probability with respect to input pixels as attribution scores (Simonyan et al., 2013). Since then, researchers have proposed numerous enhanced gradient-based methods. For example, Gradients $\times$ Input (Shrikumar et al., 2017) multiplies gradients with input values, Integrated Gradients (Sundararajan et al., 2017) accumulates gradients along a path from a baseline to the actual input, and Integrated Hessians (Janizek et al., 2021) further extends the analysis to feature interactions by computing the Hessian matrix. These methods leverage different gradient formulations to provide more accurate and stable attribution scores. A notable advancement is SmoothGrad (Smilkov et al., 2017), which generates multiple copies of the input with added Gaussian noise and computes sensitivity maps for each noisy sample. By averaging these maps, SmoothGrad reduces noise while preserving salient features that consistently influence model outputs.
@@ 5.3 Linear Approximation for Component Attribution
Given the rapid increase in model size and the combinatorial nature of searching for effective components, component attribution also employs linear approximations like LIME and Datamodels. COAR (Shah et al., 2024) attempts to decompose model behavior in terms of various model components by predicting the counterfactual impact of ablating each component, similar to many forms of causal mediation analysis. Given the computational complexity of this problem, they employ linear approximations by assigning scores to each component of a model and estimating the counterfactual effect of removing sets of components by simply summing their corresponding scores. Thus, the complexity of relationships between components is abstracted away through the linear approximation.
@@ 6 Position and Contributions
Feature, data, and component attribution methods have largely been studied as separate problems, resulting in the parallel development of similar methods from different communities with distinct terminologies. We argue that these methods can be unified into a holistic view. Having demonstrated their methodological similarities across three types of attribution, we now summarize their common concepts and challenges and identify promising research directions through cross-aspect knowledge transfer. We believe that our unified view will bridge the current fragmented landscape, make the field more accessible to newcomers, and help advance research in interpretability and beyond.
@@ 6.1 Common Concepts of Attribution Methods
As we discussed in the previous sections, attribution methods across features, data, and components can be categorized into: perturbation-based, gradient-based, and linear approximation. We provide a detailed discussion of these categories in Appendix B with Table 1 summarizing all the methods we discussed. Beyond algorithmic similarities, conceptual ideas also transfer across aspects of attribution. One example is the deliberate introduction of randomness and smoothing to enhance attribution robustness. This idea has proven effective across three attribution types through random noisy samples used by SmoothGrad for feature attribution, attribution results ensembled over multiple retrainings by TRAK for data attribution, and aggregated results from multiple corrupted runs in causal mediation analysis for component attribution. Another example is tracking and aggregating results along paths, as Integrated Gradients along paths from base input to target input in feature attribution, TracIn tracing training paths to reveal dynamic data influences in data attribution, and path patching tracking component effects along residual stream paths in component attribution. These shared concepts highlight the fundamental connections of attribution methods.
@@ 6.2 Common Challenges of Attribution Methods
Consistency Challenges refer to the variability in attribution results across multiple runs of the same method with different random seeds, making it challenging to establish stable interpretations and evaluations. This challenge is prevalent across all three types of attribution due to multiple sources of randomness in their common techniques, including sampling, learning processes with stochastic optimization, and also non-trivial hyperparameters. While some gradient-based methods can produce consistent results in a single run when they do not involve sampling or approximations for computationally intensive operations, the consistency between different gradient-based methods varies considerably, which is a problem for all three method categories and leads to the following evaluation challenges.
Evaluation Challenges arise in all three types of attribution due to multiple factors. Inconsistent results make it difficult to reliably compare methods and determine their relative accuracy. This challenge is particularly evident in gradient-based feature attribution methods. As demonstrated by Adebayo et al. (2018), they can produce contradictory attribution results and sometimes perform no better than random baselines, making them difficult to evaluate fairly. For existing evaluation metrics, counterfactual evaluation could provide more rigorous validation, but computational constraints often make this approach impractical. Task-specific evaluations offer easier alternatives but frequently lack generalizability across different contexts. Human evaluation, despite being considered the gold standard, faces scalability issues and potential biases. The diversity of evaluation metrics and their varying definitions of importance make the evaluation more challenging. An attribution result may perform well by one metric but poorly according to another. These challenges emphasize the pressing need for developing more reliable and practical attribution evaluation metrics.
6.3 Cross-Aspect Attribution Innovation
The connections among feature, data, and component attributions discussed in the sections above suggest multiple promising directions for future research. One research direction is to leverage insights from one type of attribution to develop methods for another. This can be directly identified as filling in the empty cells in Table 1. For example, while the Shapley value has been successfully applied across all three types of attribution, many other game-theoretic notions have only been used in feature attribution and not for data and component attribution. In addition, some advanced gradient techniques are common for feature and data attribution methods, but not for component attribution. The Hessian matrix, for example, has been used to obtain second-order information in Integrated Hessians for feature attribution and extensively in all IF-related methods for data attribution, and can be explored for component attribution.
@@ 6.3 Cross-Aspect Attribution Innovation
6.4 Connections to Other Areas of AI
@@ 7 Alternative Views
One alternative view suggests that feature, data, and component attribution represent fundamentally distinct research directions, each with its own terminology and research community. Despite using similar methods, these fields differ in their motivations and methodological approaches. The term “attribution” carries unique meanings in each case. Feature attribution has historically been synonymous with “explainable AI,” data attribution was part of “data-centric AI” focusing on concrete applications like data selection, and component attribution has emerged recently as one branch of “mechanistic interpretability” to reverse-engineer algorithms in foundation models. This rather independent evolution has produced distinct best practices within each domain (Guidotti et al., 2018; Hammoudeh and Lowd, 2024; Bereska and Gavves, 2024) and separate communities that may seem to resist unification. Nevertheless, we argue that meaningful unification across these three aspects is possible and will help advance interpretability research.
Researchers have also proposed various taxonomies within each type of attribution. Feature attribution methods, for example, were classified as model-agnostic or model-specific approaches (Arrieta et al., 2020). Model-agnostic methods treat models as black boxes and generate attributions based solely on input-output relationships, while model-specific methods rely on known model architectures. Although these categorizations aid method selection, they have limited generalizability to other aspects of attribution. Component attribution methods, for instance, are inherently model-specific since they analyze defined model components. Our unification approach extends beyond attribution-specific frameworks to take a broader, cross-aspect perspective. This comprehensive view uncovers connections that remain obscured when examining methods in isolation. Consider gradient-based approaches: while methods from all three types of attribution employ gradients, they compute them for different purposes with different objectives. Thus, though aspect-specific categorizations offer valuable insights, our cross-aspect framework provides a holistic understanding of the attribution landscape.
The recent work by Saphra and Wiegreffe (2024) offers another alternative view on attribution methods by examining the evolving definitions of mechanistic interpretability, a parent field of component attribution. Their study highlights how different communities within interpretability research assign distinct meanings to attributions, often guided by historical, cultural, and methodological differences, underscoring a semantic drift in how terminologies are understood. This perspective aligns with our unification position in that it recognizes attribution methods as a function of community-specific definitions and epistemic commitments, rather than strictly independent methodologies.
@@ 8 Conclusion
In this paper, we have presented a unified view on three traditionally separate attribution methods for interpretability: feature attribution, data attribution, and component attribution. While these methods have evolved independently to explain different aspects of AI systems we demonstrate that they share fundamental technical building blocks such as perturbations, gradients, and linear approximations. This unified view not only bridges the fragmented landscape of attribution methods but also identifies common challenges and opportunities for cross-attribution knowledge transfer, ultimately working toward more comprehensive approaches to advancing AI interpretability research.
9 Acknowledgements
Usha Bhalla is supported by a Kempner Institute Graduate Research Fellowship. We thank Suraj Srinivas, Yuanzhou Chen, Karim Saraipour, and Gaotang Li for their feedback on the draft of this work.
@@ Appendix B Summary of Methods
In total, thousands of attribution methods of all three types have been proposed making a comprehensive literature summary infeasible. In Table 1, we summarize the attribution methods discussed in this paper, which we believe are the representative ones and align with the unified view we presented. Most of the empty cells in Table 1 (labeled as “–”) represent methods that we believe are promising but have not yet been explored in the literature, as discussed in § 6.3. These represent research ideas that have been verified in one attribution type but remain unexplored in others. The exception is mask learning for data attribution, which we consider less promising because learning a high-dimensional mask of size $n$ jointly with the model would be infeasible when the model has not been trained.